# Bayesian Shrinkage in High-Dimensional VAR Models: A Comparative Study

### v1

@@ 4 Data Analysis
We illustrate our methods on the Canada dataset from the R package vars (Pfaff, 2008), which provides quarterly macroeconomic observations on four Canadian variables spanning $T=84$ quarters (1980Q1–2000Q4): employment (e, in log-index form), productivity (prod, in log-index form measuring labor productivity), real wages (rw, in log-index form), and the unemployment rate (U, in percent). Economic considerations suggest these variables are jointly dependent, making a Vector Auto Regression (VAR)-based approach suitable.
@@
$\widehat{\mathbf{y}}_{T+1}\;=\;\mathbf{y}_{T}\;+\;\widehat{\Delta\mathbf{y}}_{T+1},\quad\widehat{\mathbf{y}}_{T+2}\;=\;\widehat{\mathbf{y}}_{T+1}\;+\;\widehat{\Delta\mathbf{y}}_{T+2},\quad\text{etc.}$.
@@
Lag Orders from $p=1$ to $p=12$.
We fit VAR($p$) models for $p\in\{1,2,\dots,12\}$ to assess how each shrinkage method adapts to varying lag choices. Specifically, we train each VAR on the first 80 quarters of the differenced data, then produce rolling 1-step-ahead forecasts for quarters 81–84 (the holdout set). After each forecast, we roll forward by incorporating the newly observed actual data. We measure forecast accuracy in the original scale by computing both root mean squared error (RMSE), $\text{RMSE}=\sqrt{\frac{1}{4}\sum_{t=81}^{84}\left(y_{t}-\hat{y}_{t}\right)^{2}}$ and mean absolute percentage error (MAPE), $\text{MAPE}=\frac{100}{4}\sum_{t=81}^{84}\left|\frac{y_{t}-\hat{y}_{t}}{y_{t}}\right|$ over the 4 test observations.
@@ 4.1 Results
Figure 5 (left) shows the original Canadian macroeconomic variables, while the right panel displays their first-differenced versions used for estimation.
4.1.1 Aggregate Performance ($p=1,\dots,12$)
Each method’s forecast accuracy (RMSE, MAPE) for varying lag orders ($p$) is shown in Figures 6 and 7. Table 5 reports the mean and standard deviation of these metrics across all 12 fitted VAR($p$) models and all four series.
Overall, Horseshoe achieves the most consistent and accurate forecasts, attaining the lowest mean RMSE (0.51). Ridge (RMSE = 0.56) and NS (RMSE = 0.60) provide somewhat intermediate performance, while Lasso (RMSE = 0.60) and Normal (RMSE = 0.63) tend to yield slightly larger prediction errors. In terms of variability, Horseshoe’s standard deviation of RMSE (0.22) is comparable to that of Lasso and Ridge, whereas Normal exhibits the largest SD (0.26).
For MAPE, Horseshoe again stands out with an average of 0.71%, followed by NS (0.97%), Ridge (1.06%), Lasso (1.24%), and Normal (1.37%). This pattern suggests that Horseshoe effectively suppresses many small coefficients without overshrinking the larger signals, yielding robust relative‐error forecasts even as $p$ grows. By contrast, Normal’s wider prior and Lasso’s strong $\ell_{1}$ shrinkage can lead to higher MAPE in certain lag settings (see Figure 7), and NS remains susceptible to moderate forecast deterioration for larger $p$.
@@ 5 Discussion
Figure 6: Forecast performance on Canadian data with VAR($p$), $p=1,\dots,12$. Left: Out-of-sample RMSE. Right: Out-of-sample MAPE (%). Methods evaluated: Horseshoe, Lasso, ns, Normal, and Ridge. Horseshoe and Ridge consistently achieve lower RMSEs, while Lasso and Normal show moderate performance. ns experiences higher errors, especially at larger $p$.
Figure 7: Out-of-sample forecasting accuracy for each Canadian macroeconomic variable—employment ($e$), real wages ($rw$), productivity ($prod$), and unemployment ($U$)—across increasing VAR orders ($p=1,\dots,12$). The top row shows the Root Mean Squared Error (RMSE), and the bottom row shows the Mean Absolute Percentage Error (MAPE). Each colored line corresponds to one of five shrinkage methods (Horseshoe, Lasso, Normal, ns, and Ridge). Overall, Horseshoe achieves the lowest MAPE and exhibits relatively stable performance as $p$ increases. Lasso and Ridge perform well but show greater variability at higher orders.
@@
Figure 9: Out-of-sample 1 step ahead forecasts from VAR(11) and actuals for each Canadian macroeconomic variable—employment ($e$), real wages ($rw$), productivity ($prod$), and unemployment ($U$).
@@
Method Mean RMSE SD RMSE Mean MAPE (%) SD MAPE

### v2

@@ 4 Data Analysis
We illustrate our methods on the Canada dataset from the R package vars Pfaff, (2008), which provides quarterly macroeconomic observations on four Canadian variables spanning $T=84$ quarters (1980Q1–2000Q4): employment (e, in log-index form), productivity (prod, in log-index form measuring labor productivity), real wages (rw, in log-index form), and the unemployment rate (U, in percent). Economic considerations suggest these variables are jointly dependent, making a Vector Auto Regression (VAR)-based approach suitable.
@@
$\widehat{\mathbf{y}}_{T+1}\;=\;\mathbf{y}_{T}\;+\;\widehat{\Delta\mathbf{y}}_{T+1},\quad\widehat{\mathbf{y}}_{T+2}\;=\;\widehat{\mathbf{y}}_{T+1}\;+\;\widehat{\Delta\mathbf{y}}_{T+2},\quad\text{etc.}$
@@
Lag–order experiment ($p=1,\dots,12$).
For each shrinkage method we fit twelve separate VAR($p$) models ($p=1,\dots,12$). Every model is trained on the first T4T-4 quarterly differences and produces one‑step‑ahead level forecasts for 2000Q1–Q4. For a given lag order $p$ we compute
$\mathrm{RMSE}_{p}\;=\;\sqrt{\frac{1}{4}\sum_{h=1}^{4}\bigl{(}y_{T+h}-\hat{y}^{(p)}_{T+h}\bigr{)}^{2}},\qquad\mathrm{MAPE}_{p}\;=\;\frac{100}{4}\sum_{h=1}^{4}\Bigl{|}\frac{y_{T+h}-\hat{y}^{(p)}_{T+h}}{y_{T+h}}\Bigr{|}.$
Hence each method yields a set of twelve numbers $\{\mathrm{RMSE}_{p}\}_{p=1}^{12}$ and $\{\mathrm{MAPE}_{p}\}_{p=1}^{12}$. We summarize these sets by their means
$\overline{\mathrm{RMSE}}=\frac{1}{12}\sum_{p=1}^{12}\mathrm{RMSE}_{p},\quad\overline{\mathrm{MAPE}}=\frac{1}{12}\sum_{p=1}^{12}\mathrm{MAPE}_{p},$
and by their sample standard deviations
$\mathrm{SD\,RMSE}=\sqrt{\frac{1}{11}\sum_{p=1}^{12}\bigl{(}\mathrm{RMSE}_{p}-\overline{\mathrm{RMSE}}\bigr{)}^{2}},\quad\mathrm{SD\,MAPE}=\sqrt{\frac{1}{11}\sum_{p=1}^{12}\bigl{(}\mathrm{MAPE}_{p}-\overline{\mathrm{MAPE}}\bigr{)}^{2}}.$
These four statistics: $\overline{\mathrm{RMSE}}$, $\overline{\mathrm{MAPE}}$, $\mathrm{SD\,RMSE}$ and $\mathrm{SD\,MAPE}$ provide, respectively, the average forecast accuracy and the across‑lag variability reported in Table 6.
@@ 4.1 Results
4.1.1 Aggregate Performance
Overall, Horseshoe achieves the most consistent and accurate forecasts, attaining the lowest mean $\overline{\mathrm{RMSE}}$ (0.51). Ridge ($\overline{\mathrm{RMSE}}$ = 0.56) and NS ($\overline{\mathrm{RMSE}}$ = 0.60) provide somewhat intermediate performance, while Lasso (RMSE = 0.60) and Normal ($\overline{\mathrm{RMSE}}$ = 0.63) tend to yield slightly larger prediction errors. In terms of variability, Horseshoe’s standard deviation of RMSE (0.22) is comparable to that of Lasso and Ridge, whereas Normal exhibits the largest SD (0.26).
For $\overline{\mathrm{MAPE}}$, Horseshoe again stands out with an average of 0.71%, followed by NS (0.97%), Ridge (1.06%), Lasso (1.24%), and Normal (1.37%). This pattern suggests that Horseshoe effectively suppresses many small coefficients without overshrinking the larger signals, yielding robust relative‐error forecasts even as $p$ grows. By contrast, Normal’s wider prior and Lasso’s strong $\ell_{1}$ shrinkage can lead to higher MAPE in certain lag settings (see Figure 7).
@@ 4.1.2 Case Study: $p=11$
4.1.3 Practical Implications
For unemployment, the Horseshoe prior cuts the one‑step‑ahead RMSE from 0.63 to 0.51 (–19%), reducing the average error at a 6 % jobless rate from roughly 0.45 pp to 0.36 pp. System‑wide, it lowers the mean RMSE from 0.63 to 0.51 and roughly halves MAPE from 1.37 % to 0.71 % (Table 6), delivering noticeably sharper short‑term forecasts for all four Canadian macro‑series.
Relative to the strongest non‑Bayesian competitor, Ridge, Horseshoe still trims about 0.05 RMSE points and cuts MAPE by nearly one‑third. Forecast variability across lag orders remains comparable (SD RMSE 0.22 for Horseshoe vs 0.24 for Ridge; Table 6), underscoring that the accuracy gains are achieved without sacrificing stability.
@@ 5 Discussion
Conflict of Interest Disclosure
The authors declare that there are no conflicts of interest.
@@
Figure 6: One‑step‑ahead forecast errors on the Canadian hold‑out sample (2000Q1–Q4) as the VAR lag order $p$ varies. Left: root mean‑squared error $\mathrm{RMSE}_{p}$, averaged over the four macroeconomic series; Right: mean absolute percentage error $\mathrm{MAPE}_{p}$ (%) averaged over the four macroeconomic series. Lines trace the error obtained by each shrinkage estimator (Horseshoe, Ridge, Non‑parametric Shrinkage (NS), Lasso, Normal) for $p=1,\dots,12$. Horseshoe delivers the lowest and most stable errors across all lag orders. Ridge and NS perform competitively up to about $p\!\approx\!6$ but deteriorate more slowly thereafter, while Normal and Lasso show a pronounced rise in both RMSE and MAPE once p>7p\!>\!7, reflecting over‑parameterization at high lag orders.
Figure 7: Out-of-sample forecasting accuracy for each Canadian macroeconomic variable—employment ($e$), real wages (rwrw), productivity (prodprod), and unemployment (UU)—across increasing VAR orders ($p=1,\dots,12$). The top row shows the Root Mean Squared Error (RMSE), and the bottom row shows the Mean Absolute Percentage Error (MAPE). Each colored line corresponds to one of five shrinkage methods (Horseshoe, Lasso, Normal, ns, and Ridge). Overall, Horseshoe achieves the lowest MAPE and exhibits relatively stable performance as $p$ increases. Lasso and Ridge perform well but show greater variability at higher orders.
@@
Figure 9: Out-of-sample 1 step ahead forecasts from VAR(11) and actuals for each Canadian macroeconomic variable—employment ($e$), real wages (rwrw), productivity (prodprod), and unemployment ($U$).
@@
Method $\boldsymbol{\overline{\mathrm{RMSE}}}$ SD RMSE $\boldsymbol{\overline{\mathrm{MAPE}}}$ SD MAPE